OpenStudy (anonymous):

Please help Give the x and y intercept of the line 2x - 5y =10

OpenStudy (lynfran):

substitute 0 into the x... solve for y intercept substitute 0 into the y... solve for x intercept

OpenStudy (anonymous):

divide by10\[\frac{ x }{ 5 }+\frac{ y }{ -2 }=1\] x-intercept is ? y intercept is ?

OpenStudy (lynfran):

so we have \[2(0)-5y=10\]\[-5y=10\]divide by -5 we get \[y=\frac{ 10 }{ -5 }\]thats the y intercept

OpenStudy (lynfran):

so whats y=?